# A Survey on Bundle Recommendation: Methods, Applications, and Challenges

### v1

@@ 3.2.1. Backbone Models of Representation Learning
where $h_{i}^{(k)}$ represents the embedding of node $i$ in the $k$-th layer. $\mathcal{N}(i)$ represents the set of neighboring nodes of node $i$, $c_{ij}$ is the normalization coefficient, $W^{(k)}$ is the learnable weight matrix, and $\sigma$ is the non-linear activation function.
Tripartite graph: One research line (Deng et al., 2020; Lu et al., 2022; Wang et al., 2023c; Liu et al., 2020; Kim et al., 2024; Zhu et al., 2023c) focuses on constructing a user-bundle-item tripartite graph. Graph convolution operations are performed on the tripartite graph to aggregate the neighboring information into the representations of users and bundles. Among them, CL2BRec (Wang et al., 2023c) constructs a user-bundle-item tripartite graph as the global structure view to achieve augmentation at the global level. CLBR (Zhu et al., 2023c) generates a set of counterfactual views based on the original user-bundle-item graph as data augmentation, and performs propagation on the counterfactual view and the original factual graph to learn representations of users and bundles.
Hypergraph: Another line (Liu et al., 2023a; Yu et al., 2022) is to unify multiple associations between users, items, and bundles into a hypergraph, a more flexible and scalable data graph structure. HED (Liu et al., 2023a) constructs a complete hypergraph to model inter- and intra-connections among users, items, and bundles to enhance the representations. UHBR (Yu et al., 2022) updates the embeddings of users and bundles on the hypergraph, leveraging the hypergraph’s inherent ability to model higher-order associations.
Heterogeneous graph: However, the representations learned in the unified tripartite graph cannot distinguish behavioral similarities between users and content correlations between bundles from item view and bundle view, thus ignore their cooperation. Another line (Chang et al., 2020; Zhang et al., 2022, 2021b; Wang et al., 2021b; Li et al., 2023b; Du et al., 2023; Li et al., 2024b; Zou et al., 2024) is to unify users, items, and bundles into a heterogeneous graph, including a user-bundle-item interaction relational graph and a bundle-item affiliation relational graph. The representations of users and bundles are obtained by performing graph convolutional propagation and information aggregation. DT-CDBR (Zhang et al., 2021b) improves the bundle recommendation performance by using the mutual complement of the source domain and target domain on the heterogeneous graph. BGCN, CMRec, RGNN, EBRec (Chang et al., 2020; Li et al., 2023b; Wang et al., 2021b; Du et al., 2023; Zou et al., 2024) learn the representation of users and bundles by utilizing graph convolutional propagation at bundle level and item level.
Bipartite graph: A common way (Wei et al., 2023; Ke et al., 2023a, b; Ma et al., 2024a; Tan et al., 2021; Sun et al., 2022a; Ren et al., 2023; Zhao et al., 2022; Ma et al., 2022; Du et al., 2023; Jeon et al., 2024; Wu et al., 2024) is to learn the representations from bundle view and item view, which contain three bipartite graphs: namely U-B interaction graph, U-I interaction graph, and B-I affiliation graph, respectively. Many studies like (Ma et al., 2022, 2024a; Xu et al., 2023; Du et al., 2023; Jeon et al., 2024; Wu et al., 2024) employ the GNN-based recommendation framework LightGCN (He et al., 2020a) to obtain the representations of users and bundles from two views. HyperMBR, HBGCN (Ke et al., 2023a, b) use HGCN as an encoder for U-B, U-I and B-I graphs in hyperbolic space to get accurate user and bundle representations. IHBR (Tan et al., 2021) designs a node propagation based on preference transfer on both U-B and U-I graph. MIDGN (Zhao et al., 2022) applies GNNs to disentangle both U-I and B-I graph, i.e., to disentangle the user and bundle representations at the granularity of the user’s intents. CL2BRec (Wang et al., 2023c) employs a bidirectional aggregator to aggregate neighbor node information on the U-B, U-I, and B-I graphs at the local-level.
@@ 3.2.2. Representation Learning Strategy
Since there are U-I, U-B, and B-I interactions in bundle recommendation leading to two views, it is natural to focus on learning strategy between these two views when learning representations. Therefore, we classify the representation learning from the perspective of learning strategy between bundle view and item view. For ease of explanation, we illustrate the three representation learning strategies in Figure 6. One research line blindly merges the two views into a unified tripartite graph and employs graph neural networks (GNNs) or attention mechanism to aggregate the neighboring information into representations of users and bundles (Deng et al., 2020; Liu et al., 2023a; He et al., 2019b; Li et al., 2021c). We term this approach as unified learning strategy. Conversely, the second line performs representation learning upon the two views individually, and then fuses these two view-specific predictions (Chang et al., 2020; Ke et al., 2023b; Tan et al., 2021; Li et al., 2023b). This approach is defined as separate learning strategy. Finally, the third line properly models the cooperative association and encourages the mutual enhancement across the bundle and item views through techniques such as cross-view contrastive learning, metric learning, or knowledge distillation, among others (Ma et al., 2022, 2024a; Ren et al., 2023; Ke et al., 2023a). We define this as cooperative learning strategy.
[…]
Three representation learning strategies
@@
There are some existing research studies that utilized the unified strategy, which integrated users, items, and bundles into a unified user-bundle-item graph (Deng et al., 2020; Lu et al., 2022; Zhu et al., 2023c) or hypergraph (Liu et al., 2023a; Yu et al., 2022), and propagated neighboring information to get the final representations of users and bundles. However, it is obvious that user-bundle interaction density is much sparser than user-item interaction density. To confront data sparsity, counterfactual data was generated for augmentation, and the unified graph-based BR model was guided to utilize this data effectively as proposed in (Zhu et al., 2023c). Furthermore, user preference naturally propagates hierarchically from items to bundles and finally to users (He et al., 2019b; Li et al., 2020b). AttList (He et al., 2019b) applied attention networks to aggregate the item representations into a bundle representation space. Based on the aggregated bundle representations, they further aggregated them to model the latent representation of the user. HFGN (Li et al., 2020b) propagated information from lower levels to higher levels, that is, gathers information from item nodes to update bundle representations. This hierarchical approach complements the unified strategy by ensuring that detailed item-level information is effectively utilized to enhance user and bundle representations.
@@
(11) $\mathbf{e}_{u}^{B},\mathbf{e}_{b}^{B}\leftarrow U\text{-}B,$
(12) $\mathbf{e}_{u}^{I},\mathbf{e}_{b}^{I}\leftarrow U\text{-}I,B\text{-}I,$
where $\mathbf{e}_{u}^{B}$ and $\mathbf{e}_{b}^{B}$ are user embeddings and bundle embeddings learned at bundle level, $\mathbf{e}_{u}^{I}$ and $\mathbf{e}_{b}^{I}$ are user embeddings and bundle embeddings learned at item level. At each level, predictions are made independently:
(13) $\hat{y}_{B}=f_{B}(\mathbf{e}_{u}^{B},\mathbf{e}_{b}^{B}),$
(14) $\hat{y}_{I}=f_{I}(\mathbf{e}_{u}^{I},\mathbf{e}_{b}^{I}).$
@@
There are several research studies (Chang et al., 2020, 2021; Wang et al., 2021b; Zhang et al., 2021b; Ke et al., 2023b; Li et al., 2023b) etc. which propagate information at bundle level and item level, and combine bundle and item levels for final prediction. Considering that users have different preferences for different bundle categories, CMRec (Li et al., 2023b) adopts category-wise propagation at two levels, and combines category-wise user and bundle representations from both levels to estimate the likelihood of user-bundle interactions. Additionally, since the bundle-level and item-level preferences of the same user in the same time period are intertwined with each other, MULTIPLE (Li et al., 2021b) leverages the bundle-item containment relationship and obtain dual-level user preferences that incorporate information at both two levels. Specially, based on learning user preferences from bundle level and item level, BundleGT (Wei et al., 2023) further combines the information of bundling strategy into the user and bundle representations.
@@
In cooperative strategy, mutual enhancement is achieved through various techniques. One such technique is contrastive learning, which aligns and contrasts representations across the bundle view and item view to enhance representation quality. Cross-view contrastive learning has been leveraged in several research work (Ma et al., 2022; Zhao et al., 2022; Wang et al., 2023c; Du et al., 2023; Xu et al., 2023; Jeon et al., 2024). CrossCBR constructs cross-view contrastive losses to enhance the representation ability of the same user (or bundle) from different views, while decrease that of different users (or bundle). Similarly, MIDGN builds contrastive losses between global and local views. CoHEAT exploits contrastive learning to align interaction-view and affiliation-view, effectively learning latent representations. The aforementioned works is to use cross-view contrastive loss to achieve cross-view cooperation. In contrast to the works mentioned above, MultiCBR (Ma et al., 2024a) first fuses the multi-view representations into a unified one, and then adopts a simple self-supervised contrastive loss on the unified representations. Whether considering cross-views or multi-views, contrastive learning guarantees mutual enhancement of the various views. However, there exists a critical 'sampling bias' issue that similar nodes might be pushed far apart, leading to performance degradation. GPCL (Liu et al., 2023b) addresses this by devising a prototypical contrastive objective to capture the correlations between a user/bundle and its prototype. Self-supervised contrastive learning has well advanced the development of bundle recommendation. However, it may misclassify some positive samples as negative samples, resulting in suboptimal models. Inspired by this idea, DSCBR (Wu et al., 2024) integrates supervised and self-supervised contrastive learning to exploit the full potential of contrastive learning for bundle recommendation.
@@
In particular, there exists a research line, which leverages shared layers (Chen et al., 2019b; Yang et al., 2020) or shared parameters (Pathak et al., 2017) to achieve cooperation. DAM jointly models user-bundle and user-item interactions through sharing CNN layers. And BundleBPR model leverages the parameters learned through the item BPR model to estimate the preference of a user toward a bundle.
@@ 3.3. Interaction Modeling
Category Modeling Summarization References
Inner Product $\hat{y}_{ub}=\mathbf{p}_{u}^{T}\cdot aggregator(\mathbf{r}_{i})$ (Pathak et al., 2017; Li et al., 2020b)
@@
(Wang et al., 2023c; Liu et al., 2023b; Zou et al., 2024),
(Li et al., 2023b; Du et al., 2023; Li et al., 2024b) et al.
Distance Modeling Hyperbolic Distance $\hat{d}_{ub}=d_{\mathbb{H}}(\mathbf{p}^{I}_{u},\mathbf{q}^{I}_{b})+d_{\mathbb{H}}(\mathbf{p}^{B}_{u},\mathbf{q}^{B}_{b})$ (Ke et al., 2023b, a)
Neural Networks $\hat{y}_{ub}=MLP(\mathbf{p}_{u}\|\mathbf{q}_{b})$
(Chen et al., 2019b; Deng et al., 2020; He et al., 2019b; Kim et al., 2024),
(Vijaikumar et al., 2021; Agarwal et al., 2022) et al.
@@ 3.3.1. Inner Product
This approach is commonly utilized in various research works (Li et al., 2023c; Liu et al., 2023a; Yu et al., 2022; Li et al., 2021a). The bundle embeddings can be learned directly from user-bundle interactions, also can be aggregated from item embeddings using mean pooling, sum pooling or other operations (Pathak et al., 2017; Li et al., 2020b). This prediction method can be formulated as follows:
@@
However, due to the existence of bundle view and item view, some studies (Chang et al., 2020; Wang et al., 2023c; Ma et al., 2022; Zhao et al., 2022; Li et al., 2023b) learn user and bundle representations from multiple views or multiple levels. Consequently, it is essential to combine different levels of predictions for the final prediction. The most commonly used method in bundle recommendation for final prediction involves first using inner product between user and bundle representations, and then combining bundle and item levels. The abstract equation is as follows:
@@ 3.3.3. Neural Networks
Inner product models the linear interaction, and might fail to capture the complex relationships between users and bundles (He et al., 2017). Recent studies have adopted neural architectures, including Multilayer Perceptron (MLP) and Convolutional Neural Network (CNN), to mine complex and nonlinear patterns of user-bundle interactions. Neural Collaborative Filtering (NCF) is proposed to model the interaction function between each user-item pair using MLPs, achieving notable performance improvements. Expanding this approach to bundle recommendation, many studies (He et al., 2019b; Avny Brosh et al., 2022; Sun et al., 2022a; Chen et al., 2019b) etc. employ MLPs to model the user-bundle interactions. The general form is abstracted as follows:
@@
Researchers also proposed to leverage CNN based architecture for interaction modeling (Yang et al., 2020). This method first performs an outer product on user and bundle embeddings to obtain the interaction map. Above the interaction map, a stack of CNN layers is placed, which can enable the model to capture the nonlinear and higher-order correlations among users and bundles. The formula is abstracted as follows:

### v2

@@ 3.2.1. Backbone Models of Representation Learning
where $h_{i}^{(k)}$ is the embedding of node $i$ at layer $k$, $\mathcal{N}(i)$ its neighborhood, $c_{ij}$ a normalization coefficient, $W^{(k)}$ a learnable weight matrix, and $\sigma$ an activation function.
Motivated by influential surveys that categorize graph structures in recommender systems (Gao et al., 2023b; Anand and Maurya, 2025), we tailor this perspective to the task of bundle recommendation. We categorize existing studies into three representative graph types: tripartite graph, bipartite graph, hypergraph. Most works can be distinguished by whether users, bundles, and items are represented in a single graph: some unify all three types of nodes into a tripartite graph, while others decompose them into three bipartite graphs. While bipartite and tripartite graphs model pairwise relationships, a richer and more general perspective is offered by hypergraphs, which directly model the one-to-many relations within bundles. Figure 6 illustrates the three types of graph, and we summarize representative approaches for each type below.
Figure 6. Three graph types: tripartite graph, bipartite graph, hypergraph
Tripartite graph: A straightforward approach is to unify all three entity types—users, bundles, and items—into a tripartite graph. This structure allows for comprehensive modeling of all relationships in one unified view, enabling graph convolutional networks to aggregate information across different types of neighbors (Deng et al., 2020; Lu et al., 2022; Liu et al., 2020; Kim et al., 2024; Zhu et al., 2023b). Representative works include BundleNet (Deng et al., 2020), which employs a differentiable message passing framework to effectively capture the user preferences for bundles, which can incorporate the intermediate role of items between users and bundles on the user-item-bundle tripartite graph. CL2BRec (Wang et al., 2023a) constructs a user–bundle–item tripartite graph as the global structure view for data augmentation. PGAT (Lu et al., 2022) constructs a unified user–bundle–item tripartite graph and employs graph attention in neighborhood aggregation to assign differentiated weights to neighbors, enhancing the expressiveness of node embeddings.
Bipartite graph: Compared with multiple bipartite graphs, tripartite graphs are typically sparser and involve multiple heterogeneous relation types, making representation learning more challenging. In contrast, each bipartite graph contains only a single relation, leading to a denser structure and a more tractable learning process. Moreover, the separation into multiple bipartite graphs naturally aligns with the paradigm of contrastive learning, which has become an important driving force in practice. Consequently, extensive research has explored bipartite graph-based modeling for bundle recommendation (Wei et al., 2023; Ke et al., 2023a, b; Ma et al., 2024a; Tan et al., 2021; Sun et al., 2022a; Ren et al., 2023; Zhao et al., 2022; Ma et al., 2022; Du et al., 2023; Jeon et al., 2024; Wu et al., 2024), with most methods relying on three bipartite graphs: U-B, U-I, and B-I. Building on this framework, BGCN (Chang et al., 2020) unifies U–I, U–B, and B–I relations into a heterogeneous graph, where item nodes act as bridges, enabling information propagation between users and bundles allowing the learned representations to capture item-level semantics. CrossCBR (Ma et al., 2022) constructs the U–B graph as the bundle view and the U–I and B–I graphs as the item view, and then applies contrastive learning across the two views. IHBR (Tan et al., 2021) introduces a preference-transfer propagation mechanism over U–B and U–I graphs, while MIDGN (Zhao et al., 2022) disentangles user and bundle representations at the intent level by leveraging U–I and B–I graphs.
Hypergraph: Traditional graphs are limited to pairwise relations, which cannot fully capture the one-to-many nature of bundles. Hypergraphs provide a natural solution by introducing hyperedges that can directly connect a bundle to all its constituent items (Zhang et al., 2024b; Liu et al., 2024a; Yu et al., 2022, 2023; Nguyen et al., 2025). This structure allows for faithful modeling of high-order relations without decomposing them into pairwise edges. For example, HED (Liu et al., 2024a) constructs a complete hypergraph to jointly model inter- and intra-connections among users, items, and bundles, while UHBR (Yu et al., 2022) updates embeddings on the hypergraph to leverage higher-order associations. DHCL-BR (Zhang et al., 2024b) models both direct and indirect user-bundle interactions as hypergraphs and leverages dual contrastive learning to improve embedding quality. A hypergraph convolution framework is proposed for bundle representation learning, connecting all items in a bundle as hyperedges and further linking repeated items to strengthen intra-bundle correlations (Yu et al., 2023). RaMen (Nguyen et al., 2025) is a multi-strategy, multi-modal approach for bundle construction, where hypergraph message passing captures shared latent intents among item groups. These works highlight that hypergraph-based techniques effectively model complex high-order interactions.
@@ 3.2.2. Representation Learning Strategy
Since there are U-I, U-B, and B-I interactions in bundle recommendation leading to two views, it is natural to focus on learning strategy between these two views when learning representations. Therefore, we classify the representation learning from the perspective of learning strategy between bundle view and item view. For ease of explanation, we illustrate the three representation learning strategies in Figure 7. One research line blindly merges the two views into a unified tripartite graph and employs GNNs or attention mechanism to aggregate the neighboring information into representations of users and bundles (Deng et al., 2020; Liu et al., 2024a; He et al., 2019b; Li et al., 2021c). We term this approach as unified learning strategy. Conversely, the second line performs representation learning upon the two views individually, and then fuses these two view-specific predictions (Chang et al., 2020; Ke et al., 2023b; Tan et al., 2021; Li et al., 2023a). This approach is defined as separate learning strategy. Finally, the third line properly models the cooperative association and encourages the mutual enhancement across the bundle and item views through techniques such as cross-view contrastive learning, metric learning, or knowledge distillation, among others (Ma et al., 2022, 2024a; Ren et al., 2023; Ke et al., 2023a). We define this as cooperative learning strategy.
@@
There are some existing research studies that utilized the unified strategy, which integrated users, items, and bundles into a unified user-bundle-item graph (Deng et al., 2020; Lu et al., 2022; Zhu et al., 2023b) or hypergraph (Liu et al., 2024a; Yu et al., 2022), and propagated neighboring information to get the final representations of users and bundles. However, it is obvious that user-bundle interaction density is much sparser than user-item interaction density. To confront data sparsity, counterfactual data was generated for augmentation, and the unified graph-based BR model was guided to utilize this data effectively as proposed in (Zhu et al., 2023b). Furthermore, user preference naturally propagates hierarchically from items to bundles and finally to users (He et al., 2019b; Li et al., 2020b). AttList (He et al., 2019b) applied attention networks to aggregate the item representations into a bundle representation space. Based on the aggregated bundle representations, they further aggregated them to model the latent representation of the user. HFGN (Li et al., 2020b) propagated information from lower levels to higher levels, that is, gathers information from item nodes to update bundle representations. This hierarchical approach complements the unified strategy by ensuring that detailed item-level information is effectively utilized to enhance user and bundle representations.
@@
(11) $\mathbf{e}_{u}^{B},\mathbf{e}_{b}^{B}\leftarrow U\text{-}B,\mathbf{e}_{u}^{I},\;\mathbf{e}_{b}^{I}\leftarrow U\text{-}I,B\text{-}I$
where $\mathbf{e}_{u}^{B}$ and $\mathbf{e}_{b}^{B}$ are user and bundle embeddings learned at the bundle level, and $\mathbf{e}_{u}^{I}$ and $\mathbf{e}_{b}^{I}$ are user and bundle embeddings learned at the item level. Predictions at each level are made independently:
(12) $\hat{y}_{B}=f_{B}(\mathbf{e}_{u}^{B},\mathbf{e}_{b}^{B}),\;\hat{y}_{I}=f_{I}(\mathbf{e}_{u}^{I},\mathbf{e}_{b}^{I}).$
@@
There are several research studies (Chang et al., 2020, 2021; Wang et al., 2021b; Zhang et al., 2021b; Ke et al., 2023b; Li et al., 2023a) etc. which propagate information at bundle level and item level, and combine bundle and item levels for final prediction. Considering that users have different preferences for different bundle categories, CMRec (Li et al., 2023a) adopts category-wise propagation at two levels, and combines category-wise user and bundle representations from both levels to estimate the likelihood of user-bundle interactions. Additionally, since the bundle-level and item-level preferences of the same user in the same time period are intertwined with each other, MULTIPLE (Li et al., 2021b) leverages the bundle-item containment relationship and obtain dual-level user preferences that incorporate information at both two levels. Specially, based on learning user preferences from bundle level and item level, BundleGT (Wei et al., 2023) further combines the information of bundling strategy into the user and bundle representations.
@@
In cooperative strategy, mutual enhancement is achieved through various techniques. One such technique is contrastive learning, which aligns and contrasts representations across the bundle view and item view to enhance representation quality. Cross-view contrastive learning has been leveraged in several research work (Ma et al., 2022; Zhao et al., 2022; Wang et al., 2023a; Du et al., 2023; Xu et al., 2023; Jeon et al., 2024). CrossCBR constructs cross-view contrastive losses to enhance the representation ability of the same user (or bundle) from different views, while decrease that of different users (or bundle). Similarly, MIDGN builds contrastive losses between global and local views. CoHEAT exploits contrastive learning to align interaction-view and affiliation-view, effectively learning latent representations. The aforementioned works is to use cross-view contrastive loss to achieve cross-view cooperation. In contrast to the works mentioned above, MultiCBR (Ma et al., 2024a) first fuses the multi-view representations into a unified one, and then adopts a simple self-supervised contrastive loss on the unified representations. Whether considering cross-views or multi-views, contrastive learning guarantees mutual enhancement of the various views. However, there exists a critical 'sampling bias' issue that similar nodes might be pushed far apart, leading to performance degradation. GPCL (Liu et al., 2023b) addresses this by devising a prototypical contrastive objective to capture the correlations between a user/bundle and its prototype. Self-supervised contrastive learning has well advanced the development of bundle recommendation. However, it may misclassify some positive samples as negative samples, resulting in suboptimal models. Inspired by this idea, DSCBR (Wu et al., 2024) integrates supervised and self-supervised contrastive learning to exploit the full potential of contrastive learning for bundle recommendation.
@@
In particular, there exists a research line, which leverages shared layers (Chen et al., 2019a; Yang et al., 2020) or shared parameters (Pathak et al., 2017) to achieve cooperation. DAM jointly models user-bundle and user-item interactions through sharing CNN layers. And BundleBPR model leverages the parameters learned through the item BPR model to estimate the preference of a user toward a bundle.
@@ 3.3. Interaction Modeling
Category Modeling Summarization References Category Modeling Summarization References
Inner Product $\hat{y}_{ub}=\mathbf{p}_{u}^{T}\cdot aggregator(\mathbf{r}_{i})$ (Pathak et al., 2017; Li et al., 2020b) Distance Modeling
Hyperbolic Distance
$\hat{d}_{ub}=d_{\mathbb{H}}(\mathbf{p}^{I}_{u},\mathbf{q}^{I}_{b})+d_{\mathbb{H}}(\mathbf{p}^{B}_{u},\mathbf{q}^{B}_{b})$
(Ke et al., 2023b, a)
@@
Neural Networks $\hat{y}_{ub}=MLP(\mathbf{p}_{u}\|\mathbf{q}_{b})$
(Chen et al., 2019a; Deng et al., 2020; He et al., 2019b; Kim et al., 2024),
(Vijaikumar et al., 2021; Agarwal et al., 2022) et al.
@@
(Wang et al., 2023a; Liu et al., 2023b; Zou et al., 2024),
(Li et al., 2023a; Du et al., 2023; Li et al., 2024b) et al.
@@ 3.3.1. Inner Product
This approach is commonly utilized in various research works (Li et al., 2023d; Liu et al., 2024a; Yu et al., 2022; Li et al., 2021a). The bundle embeddings can be learned directly from user-bundle interactions, also can be aggregated from item embeddings using mean pooling, sum pooling or other operations (Pathak et al., 2017; Li et al., 2020b). This prediction method can be formulated as follows:
@@
However, due to the existence of bundle view and item view, some studies (Chang et al., 2020; Wang et al., 2023a; Ma et al., 2022; Zhao et al., 2022; Li et al., 2023a) learn user and bundle representations from multiple views or multiple levels. Consequently, it is essential to combine different levels of predictions for the final prediction. The most commonly used method in bundle recommendation for final prediction involves first using inner product between user and bundle representations, and then combining bundle and item levels. The abstract equation is:
@@ 3.3.3. Neural Networks
Inner product models linear interaction but might fail to capture complex relationships between users and bundles (He et al., 2017). Neural networks are widely recognized as powerful function approximators in recommender systems, capable of modeling complex nonlinear interactions (Cheng et al., 2016; Guo et al., 2017; He et al., 2017; Zhou et al., 2018). Following this trend, recent studies adopt neural architectures, including MLP and CNN, to mine nonlinear patterns of user-bundle interactions.
Neural Collaborative Filtering (NCF) is proposed to model the interaction function between each user-item pair using MLPs, achieving notable performance improvements. Expanding this approach to bundle recommendation, many studies (He et al., 2019b; Avny Brosh et al., 2022; Sun et al., 2022a; Chen et al., 2019a) etc. employ MLPs to model the user-bundle interactions. The general form is abstracted as follows:
@@
Researchers proposed leveraging CNN-based architecture for interaction modeling (Yang et al., 2020). This method first performs an outer product on user and bundle embeddings to obtain the interaction map. Above the interaction map, a stack of CNN layers is placed, enabling the model to capture the nonlinear and higher-order correlations among users and bundles. The formula is as follows: